# Transformer-based Wireless Symbol Detection Over Fading Channels

### v1

@@ V-A1 Linear Embedding
For the transmitted signal $x\in\mathbb{C}^{N_{t}}$, where each symbol is sampled from a discrete constellation set with size $C$, we represent each signal using one-hot encoding: $\tilde{x}=\text{Onehot}(x)\in\mathbb{R}^{C^{N_{t}}}.$
@@ V-A2 Decoder Layers with Masked Multi-Head Attention
Our model adopts the GPT-2 backbone with $L$ stacked decoder layers, where each layer refines the representation of the input sequence: $E^{(0)}=E,E^{(\ell+1)}=\text{DecoderLayer}(E^{(\ell)}),\ell=0,\dots,L-1.$
Each decoder layer consists of masked multi-head attention (MHA), a feedforward network (FFN), and residual connections with layer normalization. GPT-2 employs a pre-normalization scheme, where layer normalization is applied before each sub-layer [6]. The forward pass computation of the Transformer follows a standard process, which we summarize here for completeness.
Masked Multi-Head Attention (MHA)
MHA captures long-range dependencies across the input sequence, which, in our case, is established via the common channel $h_{l}$. Given the normalized input: $\tilde{E}^{(\ell)}=\text{LayerNorm}(E^{(\ell)}),$ queries ($Q$), keys ($K$), and values ($V$) are computed as: $Q=\tilde{E}^{(\ell)}W_{Q},\quad K=\tilde{E}^{(\ell)}W_{K},\quad V=\tilde{E}^{(\ell)}W_{V}.$ To enforce autoregressive behavior, a causal mask $M$ prevents attention to future tokens: $\text{MaskedAttention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+M\right)V.$ Outputs from multiple attention heads are concatenated and linearly projected: $\text{MHA}(\tilde{E}^{(\ell)})=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W_{O}.$ A residual connection is then applied: $U^{(\ell)}=E^{(\ell)}+\text{MHA}(\tilde{E}^{(\ell)}).$
Feedforward Network (FFN)
@@ V-B Model Parameters
Our specific Transformer model is designed with an embedding dimension of $d_{e}=64$, $L=8$ layers, and $h=8$ attention heads, resulting in approximately 0.42 million parameters, which is significantly smaller compared to large language models (LLMs) commonly applied in wireless communication tasks, such as those discussed by [31, 12]. For instance, even LLMs like GPT-J 6B contain over 6 billion parameters, making them approximately 14,000 times larger than our model. This compact size not only enables deployment on edge devices but also significantly shortens the inference time, enabling low-latency detection at the receiver.
V-C Training Details
In this section, we describe our data generation process and the training of the DEFINED model. Training includes a pre-training phase to equip the model with general predictive abilities and speed up convergence, followed by fine-tuning to adapt the model to scenarios with limited pilot data.
@@ V-C1 Data Generation
We generate data according to the wireless communication model described in Section III-A. Specifically, we consider both SISO and 2×2 MIMO systems and explore various modulation schemes, including BPSK, QPSK, 16QAM, and 64QAM. For each wireless system and modulation task, we generate prompts consisting of sequences with $T$ pairs, where the maximum sequence length is set to $T=31$. Both systems operate under a Rayleigh block-fading channel, with the channel coefficient sampled as $H\sim P_{H}=\mathcal{CN}(0,1)$ and remaining constant over each block. The channel noise is sampled i.i.d. from a Gaussian distribution $z_{t}\sim\mathcal{CN}(0,\sigma^{2}I)$. Within each block, the noise variance $\sigma^{2}$ is constant and shared across all noise samples. However, for different blocks, $\sigma^{2}$ is i.i.d. sampled from a uniform distribution $P_{\sigma^{2}}$ over the range $[\sigma_{\min}^{2},\sigma_{\max}^{2}]$. Varying noise levels during training enhances generalization by exposing the model to diverse channel conditions. Sampling $\sigma^{2}$ from a distribution instead of fixing it to a specific SNR improves adaptability to interference, increasing reliability in practical deployments. This variability applies only during training, while evaluation is conducted at fixed SNR levels for controlled performance assessment. The training batch size is set to 512.
@@ V-C2 ICL Pre-training
We delve into the details of model training, which is divided into two phases, as illustrated in Figure 5. First, we define ICL-training and ICL-testing as the corresponding operations on ground-truth data, represented by the clean prompt for $t=1,2,\ldots,T$: $S_{t}^{\text{ICL}}=\{y_{1},x_{1},\ldots,y_{t-1},x_{t-1},y_{t}\}$.
On the other hand, DF-training and DF-testing utilize iteratively decoded sequences with $k$ pilot data and model decision feedback, operating on the decision feedback prompt for $t=k+1,\ldots,T$: $S^{\text{DF}}_{t}=\{y_{1},x_{1},\ldots,y_{k},x_{k},y_{k+1},\hat{x}_{k+1},\ldots,y_{t-1},\hat{x}_{t-1},y_{t}\}$. Each estimation $\hat{x}_{j}$ for $j=k+1,\cdots,t-1$ is the model output, which depends on the first $k$ pilot points and preceding model decisions. This output evolves during model training and gradually converges as training progresses.
We next define the loss functions for ICL-training and DF-training. The adopted loss function is the cross-entropy loss between the model’s output and the ground-truth labels, defined as: $\text{loss}(\hat{x},x)=-\sum_{c\in\mathcal{X}}\mathbbm{1}(x=c)\log P(\hat{x}=c)$, where $\hat{x}$ is the predicted output, $x$ is the ground-truth symbol, $\mathcal{X}$ represents the modulation constellation set, and $\mathbbm{1}(\cdot)$ is the indicator function.
@@
$\mathcal{L}^{\text{DF}}(\theta)$ $=\frac{1}{N(T-k)}\sum_{i=1}^{N}\sum_{t=k+1}^{T}\text{loss}\left(f_{\theta}(S_{t,i}^{DF}),x_{t,i}\right).$ (3)
where $\theta$ represents the model parameters, and $N$ is the number of samples.
@@
In contrast, when the Transformer is trained using ICL-training but tested under DF-testing conditions, a mismatch arises: the model is trained on clean data but evaluated with noisy contexts. Despite this discrepancy, we have observed that Transformer demonstrates significant robustness to noise during DF-testing and maintains acceptable performance. Furthermore, ICL-training is approximately ten times faster than DF-training, as it eliminates the need for iterative data sampling and exclusively uses clean data.
Considering all factors, our final proposed solution is to perform ICL-training first, followed by tailored DF-training. Here, ICL-training serves as pre-training, while tailored DF-training acts as fine-tuning, similar to the pre-training and fine-tuning processes used in LLMs. Training epochs are carefully structured into two phases, as shown in Figure 5 and Figure 6. In the first phase, the model converges just before reaching a plateau, at which point we transition to the tailored DF-training method. During this transition, a spike in the training loss is observed due to the shift in the training data distribution. As shown later, ICL pre-training not only accelerates convergence but also improves recognition of clean data and ICL-testing performance.
@@ V-C3 Decision Feedback Fine-tuning
Instead of vanilla DF-training in the second phase, we employ a carefully designed fine-tuning process, as illustrated in Figure 6. The loss function is formulated as a weighted combination of the previously defined losses in Equations 2 and 3, where $\alpha$ controls the balance between them, and the hyperparameter $\alpha$ is set to 0.7 during model training,
@@
As explained in the pre-training phase, after ICL pre-training, the model is capable of general symbol detection, performing well on detection with clean data and, to some extent, on detection using the decision feedback method. Furthermore, in the fine-tuning process, the training loss is designed to emphasize decision feedback detection while retaining the model’s ability to handle clean data.
@@
VI In-Context Semi-Supervised Learning (IC-SSL)
In the previous sections, we have designed DEFINED using an approach that extends the prompt with decision pairs to improve the performance of subsequent detections. The benefits of DEFINED largely come from the pseudo-labels obtained from the decision pairs, which may not always be perfect. A natural question arises: Is there a different approach to utilize the undetected received symbols to improve detection performance?
We approach this question from the perspective of semi-supervised learning (SSL) [43]. Specifically, we propose to drop the pseudo-label and leave it null, using a semi-supervised context consisting of labeled pairs and unlabeled features. This approach allows our receiver to effectively function as an SSL estimator during inference, as illustrated in Figure 7. We hypothesize that the Transformer learns an optimal SSL estimator, enabling it to extract valuable information from unlabeled features and perform in-context semi-supervised learning during inference. This hypothesis, however, is a fundamental ML question and we leave its (dis)proof as a possible future research direction.
Figure 7: IC-SSL model architecture with $k$ pilots and $(t-k-1)$ received signals to detect $x_{t}$, $\forall t=k+1,\cdots,T$.
In this section, we first introduce the basic formulation of SSL along with its key assumptions. Then, we explain the connection to our problem formulation, introduce the new IC-SSL model, compare it to previous results, and illustrate how SSL operates within the Transformer during inference.
Semi-supervised learning (SSL) is a machine learning paradigm that lies between supervised and unsupervised learning. It leverages a small amount of labeled data alongside a large amount of unlabeled data during model training. SSL is particularly useful when the labeled data is expensive or time-consuming to acquire, while unlabeled data is abundant and easily accessible. The core idea is that unlabeled features can enhance the understanding of the feature space even without labels, thereby aiding in the design of more effective classification boundaries and ultimately improving accuracy.
The success of SSL relies on two key assumptions about the feature space:
Assumption 1 (Clustering).
Data points in the same cluster (or manifold) are more likely to share the same label.
Assumption 2 (Smoothness).
If two points are close in the space, their corresponding labels should also be close.
Figure 8: Using BPSK as an example, the received signal follows a mixture of Gaussian distributions. Relying only on labeled data, the initial decision boundary (in red) is suboptimal. By incorporating unlabeled data and assuming that the data distribution is clustered and smooth, labels can be correctly propagated to neighboring features with high probability, refining the decision boundary. This shift to the optimal boundary (in purple) demonstrates the performance improvement achieved through the inclusion of unlabeled data.
The first assumption of SSL posits that features belonging to the same class tend to cluster, and the decision boundary should pass through regions of low data density rather than intersecting the clusters. The second assumption suggests that the model’s decision function should be smooth across the data distribution, meaning that if two data points are close in the input space, their corresponding labels should also be similar. In our modulation problem, the feature space follows a mixture of Gaussian distributions, with each Gaussian centered at a constellation point and subject to Gaussian channel noise. This structure aligns with both the clustering and smoothness assumptions of SSL. As illustrated in Figure 8 using BPSK as an example, the assumptions about the received signal distribution, combined with the incorporation of unlabeled features, can effectively improve decision-making, leading to enhanced performance.
We redesign the experiment by removing the pseudo-labels and training the model using only the provided pilot pairs and unlabeled received signals. During training, the positions of the pseudo-labels are replaced with all-zero embeddings, and the same approach is applied during inference. All other training details remain the same as in the previous setup. The model, referred to as the IC-SSL model, uses an input prompt structured as for $\text{for }t=k+1,k+2,\cdots,T$: $S^{\text{SSL}}_{t}=\{y_{1},x_{1},\cdots,y_{k},x_{k},y_{k+1},\ldots,y_{t-1},y_{t}\}$, where the prompt contains $k$ pilot pairs and $(t-k)$ unlabeled features. The loss function for the IC-SSL model is defined as
$\mathcal{L}^{\text{SSL}}(\theta)=\frac{1}{N(T-k)}\sum_{i=1}^{N}\sum_{t=k+1}^{T}\text{loss}\left(f_{\theta}(S_{t,i}^{SSL}),x_{t,i}\right),$ (5)
where $f_{\theta}$ represents the Transformer model with parameters $\theta$, and the loss function is the cross-entropy loss.

### v2

@@ V-A2 Decoder Layers with Masked Multi-Head Attention
Our model adopts the GPT-2 backbone with $L$ stacked decoder layers, where each layer refines the representation of the input sequence: $E^{(0)}=E,E^{(\ell+1)}=\text{DecoderLayer}(E^{(\ell)}),\ell=0,\dots,L-1.$ Each decoder layer consists of masked multi-head attention (MHA), a feedforward network (FFN), and residual connections with layer normalization. GPT-2 employs a pre-normalization scheme, where layer normalization is applied before each sub-layer [8]. The forward pass computation of the Transformer follows a standard process, which we summarize here for completeness.
Masked Multi-Head Attention
MHA captures long-range dependencies across the input sequence, which, in our case, is established via the common channel $h_{l}$. Given the normalized input: $\tilde{E}^{(\ell)}=\text{LayerNorm}(E^{(\ell)}),$ queries ($Q$), keys ($K$), and values ($V$) are computed as: $Q=\tilde{E}^{(\ell)}W_{Q},K=\tilde{E}^{(\ell)}W_{K},V=\tilde{E}^{(\ell)}W_{V}.$ To enforce autoregressive behavior, a causal mask $M$ prevents attention to future tokens: $\text{MaskedAttention}(Q,K,V)=\text{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}+M\right)V.$ Outputs from multiple attention heads are concatenated and linearly projected: $\text{MHA}(\tilde{E}^{(\ell)})=\text{Concat}(\text{head}_{1},\dots,\text{head}_{h})W_{O}.$ A residual connection is then applied: $U^{(\ell)}=E^{(\ell)}+\text{MHA}(\tilde{E}^{(\ell)}).$
Feedforward Network
@@ V-B Model Parameters
Our specific Transformer model is designed with an embedding dimension of $d_{e}=64$, $L=8$ layers, and $h=8$ attention heads, resulting in approximately 0.42 million parameters, which is significantly smaller compared to large language models (LLMs) commonly applied in wireless communication tasks, such as those discussed by [26, 13]. For instance, even the smallest LLM used in [13], GPT-J 6B, contains over 6 billion parameters, making it approximately 14,000 times larger than DEFINED. The compact size of DEFINED not only enables its deployment on mobile devices but also significantly shortens the inference time, enabling low-latency detection at the receiver.
V-C Training Design
In this section, we describe our data generation process and the training of DEFINED.
@@ V-C1 Data Generation
We generate data according to the wireless system model described in Section III-A. Specifically, we consider both SISO and $2\times 2$ MIMO systems and explore various modulation schemes, including BPSK, QPSK, 16QAM, and 64QAM. For each wireless system and modulation task, we generate prompts consisting of sequences with $T$ pairs, where the maximum sequence length is set to $T=31$. The training data only use Rayleigh block-fading channels, with the channel coefficient sampled as $H\sim P_{H}=\mathcal{CN}(0,1)$ and remaining constant over each block. The channel noise is sampled i.i.d. from a Gaussian distribution $z_{t}\sim\mathcal{CN}(0,\sigma^{2}I)$. Within each block, the noise variance $\sigma^{2}$ is constant and shared across all noise samples. However, for different blocks, $\sigma^{2}$ is i.i.d. sampled from a uniform distribution $P_{\sigma^{2}}$ over the range $[\sigma_{\min}^{2},\sigma_{\max}^{2}]$. Varying noise levels during training enhances generalization by exposing the model to diverse channel conditions. This variability applies only during training, while evaluation is conducted at fixed SNR levels for controlled performance assessment. The training batch size is set to 512.
@@ V-C2 ICL Pre-training
We delve into the details of DEFINED pre-training using the previously described data generation, which is highly nontrivial. To prepare for the discussion, we define ICL-training and ICL-testing as the corresponding operations on ground-truth data, represented by the clean prompt for $t=1,2,\ldots,T$: $S_{t}^{\text{ICL}}=\{y_{1},x_{1},\ldots,y_{t-1},x_{t-1},y_{t}\}$. On the other hand, DF-training and DF-testing utilize iteratively decoded sequences with $k$ pilot data and model decision feedback, operating on the decision feedback prompt for $t=k+1,\ldots,T$: $S^{\text{DF}}_{t}=\{y_{1},x_{1},\ldots,y_{k},x_{k},y_{k+1},\hat{x}_{k+1},\ldots,y_{t-1},\hat{x}_{t-1},y_{t}\}$. Each estimation $\hat{x}_{j}$ for $j=k+1,\cdots,t-1$ is the model output, which depends on the first $k$ pilot points and preceding model decisions. This output evolves during model training and gradually converges as training progresses.
We next define the loss functions for ICL-training and DF-training. The adopted loss function is the cross-entropy loss between the model’s output and the ground-truth labels, defined as $\text{loss}(\hat{x},x)=-\sum_{c\in\mathcal{X}}\mathbbm{1}(x=c)\log P(\hat{x}=c)$, where $\hat{x}$ is the predicted output, $x$ is the ground-truth symbol, $\mathcal{X}$ represents the modulation constellation set, and $\mathbbm{1}(\cdot)$ is the indicator function.
@@
$\mathcal{L}^{\text{DF}}(\theta)$ $=\frac{1}{N(T-k)}\sum_{i=1}^{N}\sum_{t=k+1}^{T}\text{loss}\left(f_{\theta}(S_{t,i}^{DF}),x_{t,i}\right),$ (3)
respectively, where $\theta$ represents the model parameters, and $N$ is the number of samples.
@@
To tackle this computational challenge, one might consider directly employing ICL training for DEFINED. However, this approach would lead to a mismatch: DEFINED is trained on clean data but evaluated with noisy contexts. This discrepancy will naturally lead to performance degradation, even though we have empirically observed that ICL-training is approximately ten times faster than DF-training, as it eliminates the need for iterative data sampling and exclusively uses clean data.
Considering all factors, our final proposed solution is to perform ICL-training first, followed by tailored DF-training. Here, ICL-training serves as pre-training, while tailored DF-training acts as fine-tuning, similar to the pre-training and fine-tuning processes used in LLMs. Training epochs are carefully structured into two phases, as conceptually illustrated in Figure 5 and empirically shown in Figure 6. In the first phase, the model converges just before reaching a plateau, at which point we transition to the tailored DF-training method. During this transition, a spike in the training loss is observed due to the shift in the training data distribution. As shown later, ICL pre-training not only accelerates convergence but also improves recognition of clean data and ICL-testing performance.
@@ V-C3 Decision Feedback Fine-tuning
Instead of the vanilla DF-training in the second phase, we employ a carefully designed fine-tuning process. The new loss function is formulated as a weighted combination of the previously defined losses in Equations 2 and 3, where $\alpha$ controls the balance between them:
@@
We set hyperparameter $\alpha$ to 0.7 in the experiment.
@@
V-D Curriculum Learning
Figure 7: Training cross-entropy loss for 64QAM in SISO. The model trained without curriculum learning fails to converge, even after an extended lull period. In contrast, the model trained with curriculum learning exhibits a significantly shorter lull period, leading to faster and more stable convergence.
We further introduce curriculum learning to accelerate the training speed [20]. The core idea is to gradually increase the task difficulty, similar to a learning curriculum, progressing from simple to more challenging scenarios. In [20], this technique was applied to a linear regression problem by gradually increasing the underlying dimension and context length until the target dimension and context length were reached. The authors observed that Transformer training typically experiences an initial lull period, where the loss remains stagnant before eventually dropping sharply. Curriculum learning mitigates this lull period and accelerates convergence.
In our wireless experiments, while the channel dimension remains fixed, we progressively increase the context length every fixed number of epochs until reaching the target length. As shown in Figure 7, this approach accelerates training by reducing the lull period. Notably, for the challenging 64QAM detection task with limited pilot data and low SNR, the model may fail to converge without curriculum learning.
VI Theoretical Analysis
In this section, we present a theoretical lower bound on the detection error and analyze the generalization behavior of DEFINED under channel distribution mismatch, particularly when the SNR and Line of Sight (LoS) components differ between training and testing. Our results build upon the techniques developed in [41]. To simplify the analysis and provide cleaner insight, we focus on the BPSK SISO case, but the results can be extended to higher modulation levels. All proofs can be found in the Appendix.
To ensure a self-contained analysis, we first introduce the necessary definitions and assumptions adapted from [41]. We denote the sigmoid function by $S(x):=1/(1+\exp(-x))$. We say a data pair $(y,x)\sim\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$ if $x\in\{-1,1\}$ with equal probability and
$f(y\mid x=-1)=\mathcal{N}(\mu_{0},\Lambda),\quad f(y\mid x=1)=\mathcal{N}(\mu_{1},\Lambda),$
where $\mu_{0},\mu_{1}\in\mathbb{R}^{d}$, and $\Lambda\in\mathbb{R}^{d\times d}$ is a positive definite covariance matrix.
Assumption 1.
For an in-context test prompt $P_{\mathsf{test}}=(y_{1},x_{1},\dots,y_{k},x_{k},y_{\mathsf{query}})$, we assume: (1) $\{(y_{i},x_{i})\}_{i=1}^{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$; (2) $\mu_{0}^{\top}\Lambda^{-1}\mu_{0}=\mu_{1}^{\top}\Lambda^{-1}\mu_{1}$.
We note that Assumption 1 is satisfied for the wireless symbol detection problem. For example, in our BPSK SISO Rayleigh fading setting, the received $y_{\mathsf{query}}$ follows a distribution
$(y,x)\sim\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)=\mathcal{P}^{b}(-1,+1,\sigma^{2}I),$
where $\mu_{0}=-1$, $\mu_{1}=+1$, and $\Lambda=\sigma^{2}I$.
Following the same parameterization of the one-layer linear Transformer as in [41], the model output for a test prompt $P_{\mathsf{test}}$ with weight matrix $W$ is given by
$\widehat{x}_{\mathsf{out}}=S\left(\left(\frac{1}{k}\sum_{i=1}^{k}x_{i}y_{i}^{\top}\right)Wy_{\mathsf{query}}\right),$ (5)
The scalar output $\widehat{x}_{\mathsf{out}}\in(0,1)$ represents the model’s soft prediction. The predicted label $\widehat{x}_{\mathsf{query}}\in\{-1,1\}$ is then sampled according to: $\mathbb{P}\left(\widehat{x}_{\mathsf{query}}=1\right)=\widehat{x}_{\mathsf{out}},\mathbb{P}\left(\widehat{x}_{\mathsf{query}}=-1\right)=1-\widehat{x}_{\mathsf{out}}.$
VI-A Convergence
Following the same training procedure described in Section 3.1 of [41], and according to its Theorem 3.3, a single-layer linear Transformer with parameter $W$ trained via gradient descent and infinitely long prompts with covariance $\Lambda$ can converge to the optimal weight matrix $W^{*}=2\Lambda^{-1}$. In the matched setting where the training and test distributions share the same covariance $\Lambda$, we have the following result.
Theorem 1.
Consider a test prompt $P_{\mathsf{test}}=(y_{1},x_{1},\dots,y_{k},x_{k},y_{\mathsf{query}})$ satisfying Assumption 1. Let $\widehat{x}_{\mathsf{query}}$ denote the predicted label of the Transformer with optimal parameter $W^{*}=2\Lambda^{-1}$. Then,
$\mathbb{E}\left[\left|\mathbb{P}(\widehat{x}_{\mathsf{query}}=1\mid y_{\mathsf{query}})-\mathbb{P}(x_{\mathsf{query}}=1\mid y_{\mathsf{query}})\right|^{2}\right]$
$=$ $\frac{1}{k}\left(S^{\prime}(\mu^{\top}\Lambda^{-1}q)\right)^{2}\left[\frac{(u^{\top}\Lambda^{-1}q)^{2}}{4}+4q^{\top}\Lambda^{-1}q\right]+\;o\left(\frac{1}{k}\right),$
where $\mu=\mu_{1}-\mu_{0}$, $u=2(\mu_{1}+\mu_{0})$, and $q=y_{\mathsf{query}}$. The expectation is taken over the in-context samples $\{(y_{i},x_{i})\}_{i=1}^{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{b}(\mu_{0},\mu_{1},\Lambda)$.
Remark 1.
During testing, the prediction error of a well-trained Transformer is lower bounded at the order of $O(1/k)$. If the test prompt is also of infinite length, the Transformer can asymptotically achieve perfect prediction. Notably, Theorem 1 does not impose strong distributional assumptions on $\mu_{0}$ and $\mu_{1}$, which highlights the strong generalization ability of the trained Transformer.
Remark 2.
Since this lower bound is derived under the assumption of ground-truth in-context samples, it also serves as a lower bound for DEFINED, where the prompt consists of a limited number of pilot pairs and decision feedback (pseudo-labeled) symbols. In high-SNR scenarios, decision feedback is correct with high probability, enabling DEFINED to closely approach this theoretical bound. However, the presence of noisy feedback complicates the derivation of a tight upper bound. A rigorous theoretical analysis of Transformer-based in-context learning with decision feedback remains an open and compelling direction for future research.
VI-B Generalization
In the following, we rigorously prove that a single-layer Transformer trained at a specific SNR level can be deployed in the same Rayleigh fading channel under different SNR conditions. This result demonstrates the inherent robustness of the Transformer, which we further validate through experimental results presented in the following sections.
Theorem 2.
Suppose a test prompt $P_{\mathsf{test}}=(y_{1},x_{1},\dots,y_{k},x_{k},y_{\mathsf{query}})$ satisfies Assumption 1 with covariance $\Lambda=\sigma^{2}I$, i.e., $\{(y_{i},x_{i})\}_{i=1}^{k}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{P}^{b}(\mu_{0},\mu_{1},\sigma^{2}I)$. Now suppose the Transformer is trained on data with covariance $\Lambda^{\prime}=\xi^{2}I$, resulting in a learned optimal weight matrix $W^{*}=2\xi^{-2}I$. Then the Transformer output is $\widehat{x}_{\mathsf{out}}=S\left(\xi^{-2}\left(\frac{2}{k}\sum_{i=1}^{k}y_{i}x_{i}^{\top}\right)y_{\mathsf{query}}\right),$ and the argmax predicted label is given by $\widehat{x}_{p}=\begin{cases}1,&\text{if }\widehat{x}_{\mathsf{out}}>1/2,\\ -1,&\text{otherwise}.\end{cases}$ Then, as $k\to\infty$, we have that $\widehat{x}_{p}$ converges to the optimal predictor that maximize the accuracy between $\widehat{x}_{p}$ and $x_{\mathsf{query}}$.
Remark 3.
Theorem 2 demonstrates that a Transformer trained under one noise level ($\xi^{2}$) can generalize effectively to test-time noise levels with (1) different $\sigma^{2}$, and (2) different $\mu_{0}$, $\mu_{1}$. In a Rayleigh or Ricean fading channel, the optimal predictor that maximizes the accuracy between $\widehat{x}_{p}$ and $x_{\mathsf{query}}$ reduces to predicting based on the sign of $(\mu_{1}-\mu_{0})^{\top}x_{\mathsf{query}}$. Under SNR mismatch and LoS mismatch, only a multiplicative constant changes, while the form of the estimator remains unchanged. Furthermore, we will see in Section VII-B6 that empirical experiments corroborate these theoretical findings.